How do you solve this equation for Y?
Y-6x= -1
Please break down steps. I am having troubles with the steps. Thank you

Answer by richard1234(7193)

We can add 6x to both sides of the equation, so that the -6x on the LHS cancels.

Y -6x + 6x = -1 + 6x

Y = -1 + 6x